What is the focal width of these two parabola equations
#1 y+1=1/10(x-3)^2
#2 -1/2(y-o)^2=(x-0)

Answer by lwsshak3(11628)

What is the focal width of these two parabola equations
#1 y+1=1/10(x-3)^2
#2 -1/2(y-o)^2=(x-0)
***
convert equations to basic form:
y+1=(1/10)(x-3)^2
(x-3)^2=10(y+1)
This is an equation of a parabola that opens upwards
Its basic form: (x-h)^2=4p(y-k), (h,k)=(x,y) coordinates of the vertex
for given equation, 4p=10=focal width
..
-(1/2)(y-o)^2=(x-0)
y^2=-2x
This is an equation of a parabola that opens leftwards
Its basic form: (y-k)^2=-4p(x-h), (h,k)=(x,y) coordinates of the vertex
for given equation, 4p=2=focal width